OpenStudy (anonymous):

what is equivalent to log 50 7 rounded to three decimal places

OpenStudy (anonymous):

maybe it would be better if u use a calculator to sole it

OpenStudy (anonymous):

\[log_{50}7 =k \implies 50^k = 7\] Now take a different log of that last equation (maybe 10 or the natural log, or another one that your calculator supports.

OpenStudy (anonymous):

\[log(50^k) = log(7)\]\[\implies k\cdot log(50) = log(7)\]\[\implies k = \frac{log(7)}{log(50)}\]\[\implies log_{50}(7) = \frac{log(7)}{log(50)}\]